OpenStudy (anonymous):

How to solve cos3x=1?

OpenStudy (hysenberg):

which value of cos is 1?

OpenStudy (hysenberg):

cos0=1 so cos3x=cos0 3x=0 x=0

geerky42 (geerky42):

Not just 0. there is infinitely many values x could be, such as \(\dfrac{2}{3}\pi\), \(\dfrac{4}{3}\pi\), \(2\pi\), \(\dfrac{8}{3}\pi\) and so on